Circular No. 7110

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 990217 G. Tarei and G. Celidonio, BeppoSAX Science Operation Center, Tele-

spazio, Rome; J. in ’t Zand and H. Muller, Space Research Organization of the Netherlands, Utrecht; L. Piro and P. Soffitta, Istituto di Astrofisica Spaziale, CNR, Rome; and M. Orlandini, Istituto Tecnologie e Studio Radiazioni Extraterrestri, CNR, Bologna, report: “A γ-ray burst has been detected simultaneously by the BeppoSAX Gamma-ray Burst Monitor (GRBM) and the Wide Field Camera (WFC) no. 2 on Feb. 17.224618 UT. The GRBM lightcurve is single-peaked with a duration of ∼ 15 s and with a peak intensity of ∼ 150 counts/s in the energy band 40–700 keV. The WFC lightcurve is also single-peaked, with a peak intensity of ∼ 0.5 Crab (2–26 keV). The position derived by the WFC image is α = 3^{h}02^{m}52^{s}, δ = −53^{o}05′.6 (equinox 2000.0), with an error radius of 3′. A follow-up observation with BeppoSAX narrow-field instruments is in progress.”

SGR 1900+14 = PSR J1907+0919 Yu. P. Shitov, Pushchino Radio Astronomy Observatory, Astro Space

Center, Lebedev Physical Institute, Moscow, writes: “The soft-γ-ray repeater SGR 1900+14 was observed during 1998 Dec.–1999 Jan. using the Pushchino Bol’shaya Steerable Array (BSA) radiotelescope (area ∼ 80 000 m^{2}) and 64×20-kHz filterbank receiver at 111 MHz. In 14 observation days (BSA transit time is 200 s), we have detected the pulsed radio emission with the same period (5.161 ± 0.001 s) that was reported on IAUC 7001. The width of the integrated pulse profile at half maximum intensity is ∼ 100 ms. The estimated mean flux density at 111 MHz is ∼ 50 mJy from the best records. The dispersion-measure value was determined as 281.4 ± 0.9, which gives an estimate of the object’s distance as ∼ 6 kpc. From our timing analysis of observations made at the assumed position α = 19^{h}07^{m}14^{s}.2, δ = +9^{o}19′20′′ (equinox 2000.0; Murakami et al. 1999, Ap.J. 510, L119), we find the following best-fit parameters: barycentric period 5.161297854(83) s, period derivative dP /dt = 1.23228(34) × 10^{−10} s/s, epoch 1998 Dec. 12.460500 UT (rms timing residuals ∼ 21 ms). Thus our results show that SGR 1900+14 is identical with the radio pulsar PSR J1907+0919, representative of a new class of pulsars with a superstrong magnetic field on the order order of 8 × 10^{14} gauss, thereby confirming that this object is a ‘magnetar’ (Kouveliotou et al. 1999, Ap.J. 510, L115).”

1999 February 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green